OpenStudy (anonymous):

if a room requires 25.4 square yards of carpeting, what is the area of the floor in units of feet squared? (3 ft=1 yd)

OpenStudy (stormfire1):

Since 1 square yard = 3*3 ft: 25.4 * (3 * 3) = 228.6 sqft

OpenStudy (anonymous):

\[25.4 yd^2 \times \frac{3 ft}{1 yd} \times \frac{3 ft}{1 yd} = 228.6 ft^2\](the yards cancel)